### v1

# Balancing properties of tropical moduli maps

Abstract.
Given a family of parameterized algebraic curves over a strictly semistable pair, we show that the simultaneous tropicalization of the curves in the family forms a family of parameterized tropical curves over the skeleton of the strictly semistable pair. We show that the induced tropical moduli map satisfies a certain balancing condition, which allows us to describe properties of its image and deduce a new liftability criterion.
KC was partially supported by NSF FRG grant DMS–2053261 and a Minerva Short-Term Research Grant. XH is supported by the National Key R&\&&D Program of China 2022YFA1007100, the NSFC grant 12301057, and the ERC Consolidator Grant 770922 - BirNonArchGeom.
1. Introduction
In this paper, we study certain properties of tropicalizations of families of parameterized curves over a valued field $K$. Given a family of smooth curves $f\colon{\mathcal{C}}\to U$ together with a map ${\mathcal{C}}\to X$ to a toric variety $X$, we construct a family of tropical curves $\Gamma_{\Lambda}\to\Lambda$ together with a map $h\colon\Gamma_{\Lambda}\to{\mathbb{R}}^{n}$. This generalizes the construction for $1$-dimensional families in our previous paper [CHT23] and relies on results of de Jong [dJ96] on the existence of semistable models and of Gubler, Rabinoff, and Werner [GRW16] on skeletons of such models.
While other versions of tropicalizations of families of curves have been established in the literature, see for example [ACGS20, CCUW20, Ran22, CGM22], we propose here to focus on properties of the induced moduli map. Namely, the family $\Gamma_{\Lambda}\to\Lambda$ induces a tropical moduli map
$\alpha\colon\Lambda\to M_{g,n,\nabla}^{\mathrm{trop}}$
to the moduli space $M_{g,n,\nabla}^{\mathrm{trop}}$ of tropical parameterized curves of genus $g$, degree $\nabla$, and with $n$ marked points. As usual, it is given by sending a point in $\Lambda$ to the isomorphism class of its fiber in the tropical family. One could expect such a tropical moduli map to be a tropical map itself, i.e., to satisfy a balancing condition. Our main results give first evidence for this expectation.
Before we can state these results, recall that $M_{g,n,\nabla}^{\mathrm{trop}}$ is a (generalized) polyhedral complex with polyhedra $M_{[\Theta]}$ parameterizing tropical curves of fixed combinatorial type $\Theta$. Two classes of combinatorial types are of particular interest for our purposes:
First, weightless and $3$-valent types. The corresponding strata in $M_{g,n,\nabla}^{\mathrm{trop}}$ are maximal with respect to inclusion in the closure. Thus one may think of them as parameterizing tropically general curves.
Second, weightless and almost $3$-valent types, where we allow for a single $4$-valent vertex. A fixed weightless and almost $3$-valent stratum is contained in the closure of at most three other strata: the $3$ ways to resolve the $4$-valent vertex into a pair of adjacent $3$-valent vertices, each giving rise to a potential adjacent stratum some of which might get identified by an automorphism of the combinatorial type. Notice that the slope of the resolving edge is determined uniquely by the harmonicity condition, and each of these strata is weightless and $3$-valent.
Here is an informal version of our main result; see Theorem 3.1 for a precise statement:
Main Theorem.
Let $\alpha\colon\Lambda\to M_{g,n,\nabla}^{\mathrm{trop}}$ be as above induced by the tropicalization of a family of parameterized curves. Then,
  1. (1)
    The map $\alpha$ is (quasi-)harmonic on the interior of weightless and 3-valent strata.
  2. (2)
    The map $\alpha$ is either (quasi-)harmonic or locally combinatorially surjective on the interior of weightless and almost 3-valent strata.
More precisely, the map $\alpha$ is harmonic on codimension $1$ strata of $\Lambda$ and quasi-harmonic on strata of higher codimension. Here, quasi-harmonic means that a positive linear combination of the images of the primitive normal vectors to a stratum is contained in the image of the stratum; harmonic requires that the coefficients in the linear combination can be chosen to be $1$. Locally combinatorially surjective, on the other hand, means that the image of $\alpha$ contains points in each of the weightless and 3-valent strata adjacent to the weightless and almost 3-valent stratum. See Definitions 2.4 and 2.9.
Remark 1.1.
In [CHT22] we also considered types $\Theta$ that are not weightless or almost $3$-valent. Namely, ones that contain either a single vertex of weight $1$ or two vertices of valence $4$ adjacent to a flattened cycle. We established the corresponding liftability results for such strata in [CHT22, Lemmas 4.5 and 4.6] in the case of a $1$-dimensional family of parameterized curves. Using the arguments we present in the current article, one can also generalize these results to the setting of higher dimensional bases.
The Main Theorem allows us, in particular, to establish the following corollary, which we again state informally and refer to Corollary 3.2 for a precise statement:
Corollary.
With assumptions as in the Main Theorem we have:
  1. (1)
    If $\alpha$ is locally surjective at a point in the interior of a weightless and 3-valent stratum, it is surjective onto the closure of that stratum.
  2. (2)
    If $M_{[\Theta]}$ is a weightless and almost 3-valent stratum and $\alpha$ is surjective onto one of its adjacent weightless and 3-valent strata, it is surjective also onto all other such adjacent strata.
We view this result as a new kind of liftability/realizability criterion for parameterized tropical curves. To illustrate this, let us explain how we will use the corollary in [CHT24]. In loc. cit. we apply the corollary to the universal family ${\mathcal{C}}\to V$ over an irreducible component $V$ of a Severi variety of a toric surface $X$. Given a general point $[C]\in V$, parameterizing a curve of geometric genus $g$ on $X$, we can cut out an $N$-dimensional neighborhood of $[C]$ in $V$ by varying $N$ general point constraints, where $N$ is the dimension of $V$. Choosing the points general enough ensures that the combinatorial type $\Theta$ of the tropicalization $\mathrm{trop}(C)$ of $C$ is weightless and 3-valent, and the tropicalization of the $N$ points cuts out an $N$-dimensional neighborhood of $\mathrm{trop}(C)$ in $M_{[\Theta]}$. As it turns out, $\dim M_{[\Theta]}=N$, and hence part (1) of the corollary applies. Using part (2), it follows that the tropicalization of the universal family over $V$ contains curves in the closure of any weightless and 3-valent stratum that can be reached from $M_{[\Theta]}$ by passing through weightless and (almost) 3-valent strata.
Thus, the question of whether a given tropical curve is the tropicalization of a curve in ${\mathcal{C}}\to V$ is reduced to a purely combinatorial question, reminiscent of establishing connectivity in codimension $1$. Consequently, we show for a large class of toric surfaces that each component $V$ contains a point representing a curve with a certain tropicalization. In a second step, we show by deformation-theoretic arguments that curves with such a tropicalization can be contained in at most one irreducible component $V$, which allows us to conclude the irreducibility of Severi varieties for these toric surfaces.
2. Notation and terminology
Throughout this paper, $F$ is a complete discretely valued field with an algebraically closed residue field, and $K$ is its algebraic closure. We denote the valuation by $\nu\colon K\rightarrow\mathbb{R}\cup\{\infty\}$, the ring of integers by $K^{0}$, its maximal ideal by $K^{00}$, and the residue field by $\widetilde{K}$, and similarly for $F$.
For a subset $P\subset{\mathbb{R}}^{n}$ we denote by ${\rm Lin}(P)$ the linear subspace acting simply transitively on the affine hull of $P$. If $P$ is a polyhedron with rational slopes, then ${\rm Lin}(P)$ admits a natural integral structure induced from ${\mathbb{R}}^{n}$, i.e., the lattice of integral vectors. By an integral structure on $P$, we mean the integral structure on ${\rm Lin}(P)$.
2.1. Polyhedral complexes
Definition 2.1.
An (abstract) polyhedral complex with integral structure is a connected topological space $\Lambda$ together with a finite set $\mathcal{F}(\Lambda)$ of closed subsets, called faces of $\Lambda$, and a map $f_{W}\colon W\rightarrow N_{W}\otimes{\mathbb{R}}$ for every $W\in\mathcal{F}(\Lambda)$ such that
  1. (1)
    For every $W\in\mathcal{F}(\Lambda)$, $N_{W}$ is a lattice of finite rank;
  2. (2)
    The map $f_{W}$ is a homeomorphism from $W$ to a polyhedron with rational slopes of dimension ${\rm rank}(N_{W})$. By abuse of language, we will address $W$ itself as a polyhedron and speak about its faces, interior, dimension, integral structure, etc.;
  3. (3)
    $\Lambda=\bigsqcup_{W\in\mathcal{F}(\Lambda)}W^{\circ}$, where $W^{\circ}$ denotes the interior of $W$;
  4. (4)
    For any $W,W^{\prime}\in\mathcal{F}(\Lambda)$, the intersection $W\cap W^{\prime}$ is a union of faces of $W$ and of $W^{\prime}$. And each face in the intersection belongs to $\mathcal{F}(\Lambda)$;
  5. (5)
    If $W\in\mathcal{F}(\Lambda)$ and $W^{\prime}\in\mathcal{F}(\Lambda)$ is a face of $W$, then the integral structure on $W^{\prime}$ is the restriction of the integral structure on $W$, i.e., the linear part of $f_{W}\circ f_{W^{\prime}}^{-1}$ takes $N_{W^{\prime}}$ to a saturated sublattice in $N_{W}$.
We say that $\Lambda$ is rational if each $f_{W}(W)$ has rational vertices in $N_{W}\otimes{\mathbb{R}}$. All polyhedral complexes considered in this paper will be assumed to have integral structures, and we will call them polyhedral complexes for short.
Remark 2.2.
Axiom (2) implies that each $W\in\mathcal{F}(\Lambda)$ has a structure of a polyhedron with rational slopes, axioms (3) and (4) mean that $\Lambda$ is “glued” from polyhedra along their faces, and axiom (5) ensures that the integral structures on the polyhedra are compatible with the gluing.
Definition 2.3.
A map $f\colon\Lambda\rightarrow\Lambda^{\prime}$ is called piecewise integral affine if for each face $W\in\mathcal{F}(\Lambda)$ there is a face $W^{\prime}\in\mathcal{F}(\Lambda^{\prime})$ such that $f(W)\subseteq W^{\prime}$ and $f|_{W}\colon W\to W^{\prime}$ is integral affine.
For a face $W\in\mathcal{F}(\Lambda)$, let $\{W_{i}\}_{i}$ be the set of faces in $\mathcal{F}(\Lambda)$ containing $W$ as a codimension-one face. Then, for each $i$, the lattice $N_{W}$ is naturally identified with a saturated sublattice of $N_{W_{i}}$ of corank one. Let $\vec{e}_{i}+N_{W}$ be the generator of $N_{W_{i}}/N_{W}$ for which $f_{W_{i}}(W_{i})\subseteq f_{W_{i}}(W)+\mathbb{R}_{\geq 0}\vec{e}_{i}$, and denote by ${\rm Star}(W)$ the collection $\{\vec{e}_{i}+N_{W}\}$. We will often consider a small open neighborhood of $W^{\circ}$ in $W^{\circ}\cup\bigcup_{i}W_{i}^{\circ}$. To distinguish this set from ${\rm Star}(W)$, we denote it by ${\rm Star}(W^{\circ})$.
Definition 2.4.
Let $\Lambda$ be a polyhedral complex and $W\in\mathcal{F}(\Lambda)$ a face. A piecewise integral affine map $f\colon\Lambda\rightarrow{\mathbb{R}}^{n}$ is called quasi-harmonic at $W$ if
$\sum_{\vec{e}+N_{W}\in{\rm Star}(W)}a_{\vec{e}}\cdot\frac{\partial f}{\partial\vec{e}}\in{\rm Lin}(f(W))$
for some positive integers $a_{\vec{e}}$. It is called harmonic at $W$ if one can choose all $a_{\vec{e}}=1$. Finally, the map $f$ is called (quasi-)harmonic if it is (quasi-)harmonic at all codimension-one faces.
Remark 2.5.
Plainly, being (quasi-)harmonic at $W$ depends only on the restriction of the map $f$ to ${\rm Star}(W^{\circ})$.
2.2. Tropical curves
2.2.1. Abstract and parameterized tropical curves
The graphs in this paper are finite graphs having vertices, edges, and half-edges, called legs. We denote the set of vertices of a graph $\mathbb{G}$ by $V({\mathbb{G}})$, edges by $E({\mathbb{G}})$, and legs by $L({\mathbb{G}})$. We set $\overline{E}(\mathbb{G}):=E({\mathbb{G}})\cup L({\mathbb{G}})$. By a tropical curve $\Gamma$ we mean a finite graph $\mathbb{G}$ with ordered legs equipped with a length function $\ell\colon E(\mathbb{G})\rightarrow\mathbb{R}_{>0}$ and a weight (or genus) function $g\colon V(\mathbb{G})\rightarrow\mathbb{Z}_{\geq 0}$. We say that a curve is weightless if the weight function is identically zero. For any leg $l\in L({\mathbb{G}})$ we set $\ell(l):=\infty$.
For $e\in\overline{E}({\mathbb{G}})$, we use $\vec{e}$ to indicate a choice of orientation on $e$. If $e\in L({\mathbb{G}})$ is a leg, then it will always be oriented away from the vertex. Bounded edges will be considered with both possible orientations. We can always view tropical curves as polyhedral complexes by dividing the loops with the middle point, identifying the edges of ${\mathbb{G}}$ with bounded closed intervals of the corresponding lengths, and identifying the legs of ${\mathbb{G}}$ with semi-bounded closed intervals. This way, ${\rm Star}(v)$ is in one-to-one correspondence with the collection of oriented edges and legs having $v$ as their tail. In particular, ${\rm Star}(v)$ contains two vectors/edges for every loop based at $v$. The number of elements in ${\rm Star}(v)$ is called the valence of $v$. The genus of $\Gamma$ is defined to be $g(\Gamma)=g({\mathbb{G}}):=1-\chi(\mathbb{G})+\sum_{v\in V(\mathbb{G})}g(v)$, where $\chi({\mathbb{G}}):=b_{0}({\mathbb{G}})-b_{1}({\mathbb{G}})$ denotes the Euler characteristic of ${\mathbb{G}}$. Finally, a tropical curve $\Gamma$ is called stable if $|{\rm Star}(v)|+2g(v)\geq 3$ for every vertex $v\in V({\mathbb{G}})$.
Let $N$ be a lattice and set $N_{\mathbb{R}}:=N\otimes{\mathbb{R}}$. In the sequel, we will always let $N$ be the lattice of cocharacters of a toric variety. A parameterized tropical curve is a harmonic piecewise integral affine map $h\colon\Gamma\rightarrow N_{\mathbb{R}}$ from a tropical curve $\Gamma$ (considered as a polyhedral complex) to $N_{\mathbb{R}}$. The combinatorial type $\Theta$ of a parameterized tropical curve is defined to be the weighted underlying graph $\mathbb{G}$ with ordered legs equipped with the collection of slopes $\frac{\partial h}{\partial\vec{e}}$ for $e\in\overline{E}({\mathbb{G}})$. We denote the group of automorphisms of a combinatorial type $\Theta$ by $\mathrm{Aut}(\Theta)$, and the isomorphism class of $\Theta$ by $[\Theta]$. We define the extended degree $\overline{\nabla}$ of a parameterized tropical curve to be the sequence of slopes $\left(\frac{\partial h}{\partial\vec{l}}\right)_{l\in L({\mathbb{G}})}$, and the degree $\nabla$ the subsequence $\left(\frac{\partial h}{\partial\vec{l}_{i}}\right)$ of non-zero slopes.
2.2.2. Moduli of parameterized tropical curves
We denote by $M_{g,n,\nabla}^{\mathrm{trop}}$ the moduli space of parameterized stable tropical curves of genus $g$, degree $\nabla$, and with $n$ contracted legs. We always assume that the first $n$ legs $l_{1},\dotsc,l_{n}$ are the contracted ones. The space $M_{g,n,\nabla}^{\mathrm{trop}}$ is a generalized polyhedral complex with integral structure in the sense of [ACP15]. It is stratified by subsets $M_{[\Theta]}$ indexed by combinatorial types $\Theta$ with the fixed invariants $g,\nabla,$ and $n$. If ${\mathbb{G}}$ is the underlying graph of a combinatorial type $\Theta$, then $M_{[\Theta]}=M_{\Theta}/\mathrm{Aut}(\Theta)$, where $M_{\Theta}$ is the interior of a polyhedron $\overline{M}_{\Theta}$ in ${\mathbb{R}}^{|E({\mathbb{G}})|}\times N_{\mathbb{R}}^{|V({\mathbb{G}})|}$ and parameterizes tropical curves $h\colon\Gamma\to N_{{\mathbb{R}}}$ of type $\Theta$. The $e$-coordinate for an edge $e\in E({\mathbb{G}})$ is the length $\ell(e)$ of $e$; the $v$-coordinate for a vertex $v\in V({\mathbb{G}})$ is the coordinate of $h(v)\in N_{\mathbb{R}}$.
We call the stratum $M_{\Theta}$ weightless and $3$-valent if so is the type $\Theta$, and we call it weightless and almost $3$-valent if the type $\Theta$ is weightless and $3$-valent except for a unique $4$-valent vertex. If $M_{\Theta}$ is weightless and almost $3$-valent it is contained in the closure of at most $3$ strata, all of which are weightless and $3$-valent. Furthermore, it has codimension one in each of them.
Remark 2.6.
The nice strata (resp. simple walls) that are used in [CHT23] are particular cases of weightless and $3$-valent (resp. almost $3$-valent) strata, where the strata are in addition required to be regular, i.e., of the expected dimension.
2.2.3. Families of parameterized tropical curves over polyhedral complexes
In this subsection, we generalize the notion of one-parameter families of parameterized tropical curves in [CHT23, §3.1.3] to arbitrary dimension. Let $\Lambda$ be a polyhedral complex. Consider a datum (†) consisting of the following:
  • an extended degree $\overline{\nabla}$;
  • a combinatorial type $\Theta_{W}=\left({\mathbb{G}}_{W},\left(\frac{\partial h}{\partial\vec{\gamma}}\right)\right)$ of extended degree $\overline{\nabla}$ for each $W\in\mathcal{F}(\Lambda)$;
  • an integral affine function $\ell_{W}(\gamma,\cdot)\colon W\to{\mathbb{R}}_{\geq 0}$ for each $W\in\mathcal{F}(\Lambda)$ and $\gamma\in E({\mathbb{G}}_{W})$;
  • an integral affine function $h_{W}(u,\cdot):W\to N_{\mathbb{R}}$ for each $W\in\mathcal{F}(\Lambda)$ and $u\in V({\mathbb{G}}_{W})$;
  • a weighted contraction $\phi_{W^{\prime},W}\colon\mathbb{G}_{W^{\prime}}\to\mathbb{G}_{W}$ preserving the order of the legs for each pair of faces $W,W^{\prime}$ such that $W$ is a face of $W^{\prime}$.
For any $W\in\mathcal{F}(\Lambda)$ and $q\in W^{\circ}$, set $\Gamma_{q}:=(\mathbb{G}_{W},\ell_{q})$, where $\ell_{q}\colon E({\mathbb{G}}_{W})\to{\mathbb{R}}_{\geq 0}$ is the function defined by $\ell_{q}(\gamma):=\ell_{W}(\gamma,q)$. Denote by $h_{q}\colon\Gamma_{q}\rightarrow N_{\mathbb{R}}$ the unique piecewise affine map for which $h_{q}(u)=h_{W}(u,q)$ for all $u\in V({\mathbb{G}}_{W})$ and $\left(\frac{\partial h}{\partial\vec{l}_{i}}\right)=\overline{\nabla}$.
Definition 2.7.
Let $\Lambda$ be a polyhedral complex. We say that a datum (†) is a family of parameterized tropical curves over $\Lambda$ if the following compatibilities hold for all pairs $W,W^{\prime}\in\Lambda$ such that $W$ is a face of $W^{\prime}$, and all $q\in W^{\circ}$:
  1. (1)
    $(\Gamma_{q},h_{q})$ is a parameterized tropical curve of combinatorial type $\Theta_{W}$;
  2. (2)
    $\ell_{W}(\phi_{W^{\prime},W}(\gamma),q)=\ell_{W^{\prime}}(\gamma,q)$ for all $\gamma\in E({\mathbb{G}}_{W^{\prime}})$;
  3. (3)
    $h_{W}(\phi_{W^{\prime},W}(u),q)=h_{W^{\prime}}(u,q)$ for all $u\in V({\mathbb{G}}_{W^{\prime}})$.
A family of parameterized tropical curves over $\Lambda$ will be denoted by $h\colon\Gamma_{\Lambda}\to N_{\mathbb{R}}$ or $(\Gamma_{\Lambda},h)$. The tropical curve $(\Gamma_{q},h_{q})$ will be referred to as the fiber of $(\Gamma_{\Lambda},h)$ over $q\in\Lambda$.
Definition 2.8.
Let $\Lambda$ be a polyhedral complex, and $\alpha\colon\Lambda\to M_{g,n,\nabla}^{\mathrm{trop}}$ a continuous map. We say that $\alpha$ is piecewise integral affine if for any $W\in\Lambda$ the restriction $\alpha|_{W}$ lifts to an integral affine map $W\to\overline{M}_{\Theta}$ for some combinatorial type $\Theta$.
Let $\nabla$ be the degree associated to the extended degree $\overline{\nabla}$ by removing the zero slopes. Any family of parameterized tropical curves $h\colon\Gamma_{\Lambda}\to N_{\mathbb{R}}$ induces a piecewise integral affine map $\alpha\colon\Lambda\to M_{g,n,\nabla}^{\mathrm{trop}}$ by sending $q\in\Lambda$ to the point parameterizing the isomorphism class of the fiber $h_{q}\colon\Gamma_{q}\to N_{\mathbb{R}}$. Furthermore, $\alpha$ lifts to an integral affine map from the interior of each face of $\Lambda$ to the corresponding stratum $M_{\Theta}$. The following definition is a generalization of [CHT23, Definition 3.4].
Definition 2.9.
Let $\Lambda$ be a polyhedral complex, $\alpha\colon\Lambda\to M_{g,n,\nabla}^{\mathrm{trop}}$ a piecewise integral affine map, $W$ a face of $\Lambda$, and $\Theta$ a combinatorial type such that $\alpha(W^{\circ})\subset M_{[\Theta]}$.
  1. (1)
    Suppose $\alpha({\rm Star}(W^{\circ}))\subset M_{[\Theta]}$. We say that $\alpha$ is (quasi-)harmonic at $W$ if $\alpha_{|_{{\rm Star}(W^{\circ})}}$ lifts to a (quasi-)harmonic map ${\rm Star}(W^{\circ})\to M_{\Theta}$.
  2. (2)
    Suppose $\alpha({\rm Star}(W^{\circ}))\nsubseteq M_{[\Theta]}$. We say that $\alpha$ is locally combinatorially surjective at $W$ if for any combinatorial type $\Theta^{\prime}$ with an inclusion $\overline{M}_{\Theta}\hookrightarrow\overline{M}_{\Theta^{\prime}}$, we have $\alpha({\rm Star}(W^{\circ}))\cap M_{[\Theta^{\prime}]}\neq\emptyset$.
2.3. Families of curves
By a family of curves, we mean a flat, projective morphism of finite presentation and relative dimension one. By a collection of marked points on a family of curves, we mean a collection of disjoint sections contained in the smooth locus of the family. A family of curves with marked points is prestable if its geometric fibers have at-worst-nodal singularities. It is called (semi-)stable if so are its geometric fibers. A prestable curve with marked points defined over a field is called split if the irreducible components of its normalization are geometrically irreducible and smooth and the preimages of the nodes in the normalization are defined over the ground field. A family of prestable curves with marked points is called split if all of its fibers are so; cf. [dJ96, § 2.22]. If $U\subset Z$ is open, and $({\mathcal{C}},\sigma_{\bullet})$ is a family of curves with marked points over $U$, then by a model of $({\mathcal{C}},\sigma_{\bullet})$ over $Z$ we mean a family of curves with marked points over $Z$, whose restriction to $U$ is $({\mathcal{C}},\sigma_{\bullet})$; and similarly, for prestable and semi-stable models.
2.4. Toric varieties and parameterized curves
For a toric variety $X$, we denote the lattice of characters by $M$, of cocharacters by $N$, and the monomial functions by $x^{m}$ for $m\in M$. A parameterized curve in a toric variety $X$ is a smooth projective curve with marked points $(C,\sigma_{\bullet})$ together with a map $f\colon C\to X$ such that $f(C)$ does not intersect orbits of codimension greater than one, and the image of $C\setminus\left(\bigcup_{i}\sigma_{i}\right)$ under $f$ is contained in the dense orbit of $X$.
A family of parameterized curves $f\colon{\mathcal{C}}\rightarrow X$ over $K$ consists of the following data:
  1. (1)
    a family of smooth marked curves $({\mathcal{C}}\to B,\sigma_{\bullet})$ over a base $B$, and
  2. (2)
    a map $f\colon{\mathcal{C}}\rightarrow X$, such that for any geometric point $p\in B$ the restriction ${\mathcal{C}}_{p}\to X$ is a parameterized curve.
2.5. Tropicalization of parameterized curves
Let $f\colon C\to X$ be a parameterized curve, and $C^{0}\to{\rm Spec}(K^{0})$ a prestable model. Denote by $\widetilde{C}$ the fiber of $C^{0}$ over the closed point of ${\rm Spec}(K^{0})$. The tropicalization $\mathrm{trop}(C)$ of $C$ with respect to the model $C^{0}$ is the tropical curve $\Gamma=(\mathbb{G},\ell)$ defined as follows: The underlying graph ${\mathbb{G}}$ is the dual graph of the central fiber $\widetilde{C}$, i.e., the vertices of $\mathbb{G}$ correspond to irreducible components of $\widetilde{C}$, the edges – to nodes, the legs – to marked points, and the natural incidence relation holds. For a vertex $v$ of $\mathbb{G}$, its weight is defined to be the geometric genus of the corresponding component $\widetilde{C}_{v}$ of the reduction $\widetilde{C}$. As for the length function, if $e\in E({\mathbb{G}})$ is the edge corresponding to a node $z\in\widetilde{C}$, then $\ell(e)$ is defined to be the valuation of $\lambda$, where $\lambda\in K^{00}$ is such that étale locally at $z$, the total space of $C^{0}$ is given by $xy=\lambda$. Although $\lambda$ depends on the étale neighborhood, its valuation does not, and hence the length function is well-defined. Finally, notice that the order on the set of marked points induces an order on the set of legs of $\Gamma$.
Next, we explain how to construct the parameterization $h\colon\mathrm{trop}(C)\to N_{\mathbb{R}}$. Let $\widetilde{C}_{v}$ be an irreducible component of $\widetilde{C}$. Then, for any $m\in M$, the pullback $f^{*}(x^{m})$ of the monomial $x^{m}$ is a non-zero rational function on $C^{0}$ since the preimage of the dense orbit is dense in $C$. Thus, there is a $\lambda_{m}\in K^{\times}$, unique up to an element invertible in $K^{0}$, such that $\lambda_{m}f^{*}(x^{m})$ is an invertible function at the generic point of $\widetilde{C}_{v}$. The function $h(v)$, associating to $m\in M$ the valuation $\nu(\lambda_{m})$, is clearly linear, and hence $h(v)\in N_{\mathbb{R}}$. The parameterization $h\colon\mathrm{trop}(C)\to N_{\mathbb{R}}$ is defined to be the unique piecewise integral affine function with values $h(v)$ at the vertices of $\mathrm{trop}(C)$, whose slopes along the legs satisfy the following: for any leg $l$ and $m\in M$ we have $\frac{\partial h}{\partial\vec{l}}(m)=-\mathrm{ord}_{\sigma_{i}}f^{*}(x^{m})$, where $\sigma_{i}$ is the marked point corresponding to $l$. Then $h\colon\mathrm{trop}(C)\to N_{\mathbb{R}}$ is a parameterized tropical curve, by [Tyo12, Lemma 2.23]. It is called the tropicalization of $f\colon C\to X$ with respect to the model $C^{0}$.
Plainly, the tropical curve $\mathrm{trop}(C)$ is independent of the parameterization and depends only on $C^{0}$. If the family $C\to X$ is stable and $C^{0}$ is the stable model, then the corresponding tropicalization is called simply the tropicalization of $C$ (resp. of $f\colon C\to X$).
Remark 2.10.
Let $l$ be the leg corresponding to a marked point $\sigma_{i}$. If $\sigma_{i}$ is mapped to the boundary divisor $D$, then, by definition, $f(\sigma_{i})$ belongs to a unique component of $D$, and in particular, to the smooth locus of $D$. Thus, the direction of the slope $\frac{\partial h}{\partial\vec{l}}$ is completely determined by the irreducible component of $D$ containing the image of $\sigma_{i}$. Furthermore, the integral length of $\frac{\partial h}{\partial\vec{l}}$ is the multiplicity of $\sigma_{i}$ in $f^{*}D$. Note also that $\frac{\partial h}{\partial\vec{l}}=0$ if and only if $f(\sigma_{i})$ is contained in the dense orbit of $X$.
2.6. Strictly semistable pairs and their tropicalization
To tropicalize families, we will first tropicalize the base of the family. To do so, we shall use higher-dimensional tropicalizations. Different versions of such tropicalizations have appeared in the literature; see, for example, [MN15, ACGS20, CCUW20, Ran22, CGM22]. Here we will follow the construction by Gubler, Rabinoff, and Werner [GRW16]. Let us recall their construction.
Definition 2.11.
([GRW16, Definition 3.1]) A strictly semistable pair $(B^{0},H^{0})$ over $K^{0}$ consists of an irreducible proper flat scheme $B^{0}$ over $K^{0}$ and a sum $H^{0}=\sum H_{k}$ of distinct effective Cartier divisors $H_{k}$ on $B^{0}$ such that $B^{0}$ is covered by open subsets $U$ which admit an étale morphism
$\varphi\colon U\rightarrow\mathrm{Spec}(K^{0}[x_{0},\dotsc,x_{d}]/\langle x_{0}\cdot\ldots\cdot x_{a}-\lambda\rangle)$ (2.1)
for some $a\leq d$ and $\lambda\in K^{00}$. Moreover, we assume that each $H_{k}$ has irreducible support and $H_{k}\cap U$ is either empty or defined by $\varphi^{*}(x_{j})$ for some $a+1\leq j\leq d$. Similarly, define a strictly semistable pair over $F^{0}$ by replacing $K^{0}$ with $F^{0}$.
For a strictly semistable pair $(B^{0},H^{0})$, denote by $\widetilde{B}$ the special fiber of $B^{0}$ with irreducible components $\widetilde{B}_{j}$. Set
$D\coloneqq\widetilde{B}\cup(\bigcup_{k}H_{k})\subset B^{0}$ (2.2)
with irreducible components $\{D_{i}\}_{i\in I}=\{H_{k}\}_{k}\cup\{\widetilde{B}_{j}\}_{j}$. Then the $\widetilde{B}_{j}$’s are called the vertical components of $D$ and the $H_{k}$’s the horizontal components. The divisor $D$ admits a stratification by the irreducible components of $\bigcap_{i\in J}D_{i}\backslash\big{(}\bigcup_{i\not\in J}D_{i}\big{)}$ for any $J\subseteq I$. There is a natural partial order on the set of strata of $D$ given by $S\leq T$ if and only if $S\subseteq\overline{T}$. Let $\mathrm{Str}(B^{0},H^{0})$ denote the set of strata with vertical support, namely, those contained in $\widetilde{B}$. Then for any $S\in\mathrm{Str}(B^{0},H^{0})$, the closure $\overline{S}$ is smooth over $\widetilde{K}$, cf. [GRW16, §3.15]. In particular, each component of the reduction $\widetilde{B}_{j}$ is smooth over $\widetilde{K}$. Notice that our notation differs slightly from the one in loc. cit., where the set of strata $\mathrm{Str}(B^{0},H^{0})$ is denoted by ${\rm str}(\widetilde{B},H^{0})$.
We can now describe the skeleton of a strictly semistable pair $(B^{0},H^{0})$ following [GRW16, §4]. Notice, however, that in loc. cit. the construction is described for a formal strictly semistable pair obtained from $(B^{0},H^{0})$ by a formal completion with respect to a non-zero element of $K^{00}$. To a stratum $S\in\mathrm{Str}(B^{0},H^{0})$, one associates a polyhedron $\Delta_{S}$ as follows. Pick a sufficiently small open neighborhood $U$ of a general point of $S$ such that there exists an étale morphism $\varphi$ as in (2.1) and such that for any stratum $T$ the following holds: $S\leq T$ if and only if $T\cap U\neq\emptyset$. Let $\{H_{i}\}_{1\leq i\leq b}$ be the horizontal divisors that intersect $U$, where $0\leq b\leq d-a$. We may assume that $H_{i}$ is defined by $\varphi^{*}(x_{a+i})$, and that $S$ is given by $\varphi^{*}(x_{0})=\cdots=\varphi^{*}(x_{a+b})=0$. Set $\Delta_{S}:=\Delta(a,\lambda)\times\mathbb{R}^{b}_{\geq 0}$, where
$\Delta(a,\lambda)=\{y=(y_{0},\dotsc,y_{a})\in\mathbb{R}_{\geq 0}^{a+1}|y_{0}+\cdots+y_{a}=\nu(\lambda)\},$
and notice that it is strictly convex, i.e., contains no lines. Following [GRW16], we call $\nu(\lambda)$ the length of $S$.
For $U$ and $\Delta_{S}$ as above, consider the locus $U_{S}\subset U\setminus D$ of $K$-points $p$, whose specialization belongs to the union of strata $\cup_{S\leq T}T$. There is a natural tropicalization map $\tau_{U}\colon U_{S}\rightarrow\Delta_{S}$ given by
$p\mapsto\left(\nu\left(\varphi^{*}(x_{0})(p)\right),\dotsc,\nu\left(\varphi^{*}(x_{a+b})(p)\right)\right).$
Then $\Delta_{S}$ is independent of the choice of $U$ and $\varphi$, up to reordering the coordinates, and the tropicalization map is also independent of the choice of $\varphi$, and compatible on the intersections for different choices of $U$; cf. [GRW16, §4.3].
If $S\leq T$ are strata, then up to reordering of the coordinates, there exists $a^{\prime}\leq a$ and $b^{\prime}\leq b$, such that $T$ is given locally by $\varphi^{*}(x_{0})=\cdots=\varphi^{*}(x_{a^{\prime}})=\varphi^{*}(x_{a+1})=\cdots=\varphi^{*}(x_{a+b^{\prime}})=0.$ Therefore, $\Delta_{T}$ can be identified naturally with a face of $\Delta_{S}$. Moreover, the tropicalization maps agree on $U_{T}$. Since $B^{0}$ is proper over ${\rm Spec}(K^{0})$, any $K$-point specializes to one of the strata in $\mathrm{Str}(B^{0},H^{0})$, and therefore, the tropicalization maps glue to a well-defined tropicalization map $\tau$ from the set of $K$-points of $B^{0}\setminus D$ to the limit $S(B^{0},H^{0}):=\bigcup_{S\in\mathrm{Str}(B^{0},H^{0})}\Delta_{S}$; cf. [GRW16, §4.6, §4.9]. In the sequel, e.g., in Theorem 3.1, we will abuse the notation and denote the tropicalization map $\tau$ by $\mathrm{trop}$.
Definition 2.12.
The polyhedral complex $S(B^{0},H^{0}):=\bigcup_{S\in\mathrm{Str}(B^{0},H^{0})}\Delta_{S}$ is called the skeleton of the strictly semistable pair $(B^{0},H^{0})$.
The skeleton is naturally a rational polyhedral complex with an integral structure. Indeed, any face $\Delta_{S}$ admits a natural embedding $\Delta_{S}=\Delta(a,\lambda)\times\mathbb{R}^{b}_{\geq 0}\rightarrow{\mathbb{R}}^{a+1+b}$, which induces the integral structure on $\Delta_{S}$. By construction, these integral structures are compatible for different faces of the skeleton.
3. Main theorem
In this section, we fix a pair of dual lattices $M$ and $N$, and a toric variety $X$ with lattice of monomials $M$. Let ¯$\overline{{\mathcal{M}}}_{g,n}$ be the algebraic stack over $\mathbb{Z}$ classifying stable $n$-pointed curves of genus $g$. Then ¯$\overline{{\mathcal{M}}}_{g,n}$ admits a finite surjective morphism from a projective scheme $\overline{M}$, over which there is a universal family of curves, cf. [dJ96, §2.24].
Theorem 3.1.
Let $(B^{0},H^{0})$ be a strictly semistable pair over $K^{0}$ and $({\mathcal{C}}^{0}\rightarrow B^{0},\sigma^{0}_{\bullet})$ a split family of stable marked curves over $B^{0}$. Let $B$, $H$ and ${\mathcal{C}}$ be the generic fibers of $B^{0}$, $H^{0}$, and ${\mathcal{C}}^{0}$, respectively. Suppose ${\mathcal{C}}^{0}$ is obtained as a pullback of the universal family over $\overline{M}$ and is smooth over $B^{\prime}:=B\backslash H$. Suppose furthermore that $f\colon{\mathcal{C}}|_{B^{\prime}}\rightarrow X$ is a family of parameterized curves over $B^{\prime}$. Then there exists a family of parameterized tropical curves $h\colon\Gamma_{\Lambda}\to N_{\mathbb{R}}$ over $\Lambda:=S(B^{0},H^{0})$ such that
  1. (1)
    For any $\eta\in B^{\prime}(K)$, the fiber of $(\Gamma_{\Lambda},h)$ over $\mathrm{trop}(\eta)$ is the tropicalization of $f\colon{\mathcal{C}}_{\eta}\to X$;
  2. (2)
    For any face $W\in{\mathcal{F}}(\Lambda)$ of codimension 1 (resp. of codimension at least 2) such that the tropical curves over the interior of $W$ are weightless and $3$-valent except for at most one $4$-valent vertex, the induced map $\alpha\colon\Lambda\to M^{\mathrm{trop}}_{g,n,\nabla}$ is either harmonic (resp. quasi-harmonic) or locally combinatorially surjective along $W$.
The following corollary is useful in applications, see, e.g., [CHT24].
Corollary 3.2.
Let $X$ be a toric variety and $f\colon{\mathcal{C}}\rightarrow X$ a family of parameterized curves defined over the field $K$. Assume that the base $B^{\prime}$ of the family is quasi-projective, and consider the set of tropicalizations
$\Sigma:=\left\{\left[\mathrm{trop}(f_{\eta},C_{\eta})\right]\;:\;\eta\in B^{\prime}(K)\right\}\subseteq M^{\mathrm{trop}}_{g,n,\nabla}.$
Let $M_{[\Theta]}\subset M^{\mathrm{trop}}_{g,n,\nabla}$ be a stratum. Then,
  1. (1)
    The closure $\overline{\Sigma}\subseteq M^{\mathrm{trop}}_{g,n,\nabla}$ is the image of a rational polyhedral complex $\Lambda$ under a piecewise integral affine map;
  2. (2)
    If $M_{[\Theta]}$ is weightless and $3$-valent and $\dim(\Sigma\cap M_{[\Theta]})=\dim M_{[\Theta]}$, then $\overline{M}_{[\Theta]}\subseteq\overline{\Sigma}$;
  3. (3)
    If $M_{[\Theta]}$ is weightless and almost $3$-valent and $\overline{\Sigma}$ contains one of its adjoint weightless and $3$-valent strata, then $\overline{\Sigma}$ contains all other such strata as well.
Proof.
After replacing $B^{\prime}$ with a finite covering, we may assume that the family of marked curves over $B^{\prime}$ is a pullback of the universal family over the projective scheme $\overline{M}$. Plainly, it suffices to prove the corollary for each irreducible component of $B^{\prime}$. Furthermore, we may replace $B^{\prime}$ with $B^{\prime}_{\rm red}$, and hence we will assume that $B^{\prime}$ is integral, i.e., a quasi-projective variety. Thus, there is a projective variety $B$ containing $B^{\prime}$ as a dense open subset. After replacing $B$ with the closure of the graph $B^{\prime}\to\overline{M}\times B$, we may assume that the family of marked curves over $B^{\prime}$ extends to $B$, and the latter family is also a pullback of the universal family over $\overline{M}$. Set $H:=B\setminus B^{\prime}$.
By [Sta20, Tag 0C2S], $B$ admits a flat projective model $B^{0}$ over $K^{0}$. Let $\overline{M}^{0}$ be the trivial model of $\overline{M}$ over $K^{0}$. After replacing $B^{0}$ with the schematic image of the graph $B\to\overline{M}^{0}\times_{{\rm Spec}(K^{0})}B^{0}$, which is also flat over $K^{0}$ by loc. cit., we may assume that the family of marked curves extends to $B^{0}$. By [dJ96, Lemma 5.3], there exist a projective alteration $\varphi_{1}\colon B_{1}^{0}\to B^{0}$ such that the pullback of ${\mathcal{C}}$ to $B_{1}^{0}$ is split over $B_{1}^{0}$. Replacing $B_{1}^{0}$ with an irreducible component that dominates $B^{0}$ if necessary, we may assume that $B_{1}^{0}$ is integral. Since the loci $\Sigma$ associated to $B^{\prime}$ and to $\varphi_{1}^{-1}(B^{\prime})$ coincide, we may assume that $B_{1}^{0}=B^{0}$ and hence the family ${\mathcal{C}}$ is split over $B^{0}$. Finally, by [dJ96, Theorem 6.5], there exists an alteration $\varphi_{2}\colon B_{2}^{0}\to B^{0}$ such that $B_{2}^{0}$ is projective over $K^{0}$, the generic fiber of $B_{2}^{0}\to{\rm Spec}(K^{0})$ is irreducible, and the pair $\left(B_{2}^{0},\overline{\varphi_{2}^{-1}(H)}_{\rm red}\right)$ is strictly semistable. Here $\overline{\varphi_{2}^{-1}(H)}_{\rm red}$ is the closure of $\varphi_{2}^{-1}(H)_{\rm red}$ in $B_{2}^{0}$. As before, we may assume that $B_{2}^{0}=B^{0}$.
Denote by $H^{0}$ the union of the horizontal components of $H\cup\tilde{B}$. Then $(B^{0},H^{0})$ is a strictly semistable pair, and the family of marked curves extends to $B^{0}$ by pulling back the universal family over $\overline{M}$. Therefore, we may assume that we are given all the data $B^{0},H^{0},{\mathcal{C}}^{0},\sigma^{0}_{\bullet},B^{\prime}$ satisfying the assumptions of Theorem 3.1. Consider the tropicalization $h\colon\Gamma_{\Lambda}\to N_{\mathbb{R}}$ and the induced map $\alpha\colon\Lambda\to M^{\mathrm{trop}}_{g,n,\nabla}$ provided by Theorem 3.1.
(1) Since $\alpha$ is piecewise integral affine, it follows that $\alpha(\Lambda)\subseteq M^{\mathrm{trop}}_{g,n,\nabla}$ is closed. Now, by Theorem 3.1 (1), $\Sigma\subset\alpha(\Lambda)$, and therefore $\overline{\Sigma}\subset\alpha(\Lambda)$. Furthermore, $\alpha^{-1}(\Sigma)$ is dense in $\Lambda$. Thus, $\Sigma$ is dense in $\alpha(\Lambda)$, and hence $\overline{\Sigma}=\alpha(\Lambda)$ as asserted.
(2) Set $k:=\dim M_{[\Theta]}$, and assume to the contrary that $\overline{M}_{[\Theta]}\nsubseteq\alpha(\Lambda)$. By (1), $\alpha(\Lambda)\subseteq M^{\mathrm{trop}}_{g,n,\nabla}$ is closed, and therefore $M_{[\Theta]}\nsubseteq\alpha(\Lambda)$. Furthermore, there is a face $W\in{\mathcal{F}}(\Lambda)$ such that $\alpha(W^{\circ})$ has dimension $k-1$, $\alpha\left({\rm Star}(W^{\circ})\right)$ has dimension $k$, and $\alpha(W^{\circ})$ belongs to the relative boundary of $\alpha(\Lambda)$ in $M_{[\Theta]}$.
After replacing $W$ with one of its faces, we may assume that $\dim(W)=k-1$. Indeed, since $\dim(\alpha(W^{\circ}))=k-1$, the dimension of $W$ is at least $k-1$. If it is greater than $k-1$, then the kernel of the differential $d\alpha$ along $W$ is non-trivial. But $\Lambda$ is the skeleton of a strictly semistable pair, and therefore $W$ contains no lines. Thus, $\alpha(W)=\alpha(\partial W)$, and by induction, we can find a face of $W$ of dimension $k-1$ satisfying the requirements.
By construction, $\alpha\left({\rm Star}(W^{\circ})\right)$ belongs to a half-space supported by $\alpha(W^{\circ})$ and the differential $d\alpha$ is not identically zero on ${\rm Star}(W^{\circ})$. Therefore $\alpha$ is not quasi-harmonic along $W$ contradicting Theorem 3.1 (2).
(3) Let $M_{[\Theta^{\prime}]}$ be a weightless and $3$-valent stratum adjacent to $M_{[\Theta]}$ that is contained in $\overline{\Sigma}$, and $M_{[\Theta^{\prime\prime}]}$ be another weightless and $3$-valent stratum adjacent to $M_{[\Theta]}$. Then there exists a face $W^{\prime}\in\mathcal{F}(\Lambda)$ and a face $W$ of $W^{\prime}$ of codimension one such that $\alpha\left((W^{\prime})^{\circ}\right)\subseteq M_{[\Theta^{\prime}]}$ has dimension $\dim M_{[\Theta^{\prime}]}$, and $\alpha(W^{\circ})\subseteq M_{[\Theta]}$. Thus, $\alpha$ is locally combinatorially surjective along $W$ by Theorem 3.1 (3). In particular, there exists a face $W^{\prime\prime}\in\mathcal{F}(\Lambda)$ containing $W$ as a codimension-one face such that $\alpha\left((W^{\prime\prime})^{\circ}\right)\subseteq M_{[\Theta^{\prime\prime}]}$, and hence
$\dim\alpha\left(W^{\prime\prime}\right)\geq\dim\alpha(W)+1=\dim\left(M_{[\Theta^{\prime\prime}]}\right)$
since $\alpha|_{W^{\prime\prime}}$ is an affine map and $\alpha(W)\subsetneq\alpha(W^{\prime\prime})$. Therefore, $\overline{M}_{[\Theta^{\prime\prime}]}\subseteq\overline{\Sigma}$ by (2), and we are done. ∎
Remark 3.3.
The above results, while applicable to the more general case of weightless and (almost) $3$-valent strata, are conceived with a view towards their application in [CHT24]. There we will use simple walls and nice strata of [CHT23], which are weightless and (almost) $3$-valent strata, which in addition are regular, i.e., of the expected dimension. Two weightless and (almost) $3$-valent, but not regular, strata were studied already in [CHT22]:
A non-regular, weightless and almost $3$-valent type $\Theta$ is given by a contracted loop adjacent to the $4$-valent vertex, and no other contracted edges. In this case, the stratum $M_{[\Theta]}$ has dimension one more than expected. While Corollary 3.2 (2) does not apply in this case, since the stratum is not $3$-valent, the claim still holds: since $\alpha$ is not locally combinatorially surjective at $M_{[\Theta]}$ by the proof of [CHT22, Lemma 4.6], it is (quasi-)harmonic by Theorem 3.1. This in turn is the only statement we used to deduce Corollary 3.2 (2).
A non-regular, weightless and $3$-valent type $\Theta$ on the other hand appears when $\Theta$ is $3$-valent but contains a single flattened cycle as in [CHT22, Definition 3.7]. In this case, we again have dimension one more than expected, and $M_{[\Theta]}$ contains curves that are not realizable [CHT22, Proposition 3.8]. In particular, the condition of local surjectivity of $\alpha$ in Corollary 3.2 (2) can never be satisfied. However, one can describe the realizable sublocus in $M_{[\Theta]}$ explicitly, and (quasi-)harmonicity of $\alpha$ again ensures that if $\alpha$ is locally surjective at a point in this locus, it is surjective onto the whole realizable locus in $M_{[\Theta]}$.
The rest of the section is devoted to the proof of Theorem 3.1. Since $K$ is the algebraic closure of a complete discretely valued field $F$, any $K$-scheme of finite type is defined over a finite extension $F^{\prime}$ of $F$, which is also a complete discretely valued field since so is $F$. Thus, we may view a $K$-scheme of finite type as the base change of a scheme over a finite extension of $F$. In the proofs below, we will work over $F^{\prime}$ in order to have a well-behaved total space of the families we consider. To ease the notation, we will assume that all points we are interested in are defined already over $F$ and integral models – over $F^{0}$. For the rest of the proof, we fix a uniformizer $\pi\in F^{00}$, and assume without loss of generality that $\nu(\pi)=1$. In particular, we may assume that $\eta$ and $f$ are defined over $F$ and $({\mathcal{C}}^{0}\rightarrow B^{0},\sigma^{0}_{\bullet})$ over $F^{0}$.
Throughout the proof, we will use the following notation: $\widetilde{B}$ and $\widetilde{\mathcal{C}}$ will denote the special fibers, $s:=\mathrm{red}(\eta)\in\widetilde{B}$ the reduction of $\eta$, and $\widetilde{\mathcal{C}}_{s}$ the corresponding fiber of ${\mathcal{C}}^{0}\to B^{0}$. We set $D_{B^{0}}:=\widetilde{B}\cup H$ and $D_{{\mathcal{C}}^{0}}:=\widetilde{\mathcal{C}}\cup{\mathcal{C}}_{H}\cup\left(\bigcup_{j}\sigma_{j}^{0}\right)$, where ${\mathcal{C}}_{H}$ is the restriction of ${\mathcal{C}}$ to $H$. Then, the pullback of any monomial function $f^{*}(x^{m})$ is regular and invertible on ${\mathcal{C}}^{0}\setminus D_{{\mathcal{C}}^{0}}$ by the assumptions of the theorem. Recall that $D_{B^{0}}$ has a stratification, and we denote by $\mathrm{Str}(B^{0},H^{0})$ the set of strata with vertical support.
Let $S\in\mathrm{Str}(B^{0},H^{0})$ be the stratum containing $s$, and $s\in U\subset B^{0}$ an open neighborhood admitting an étale morphism $\varphi\colon U\rightarrow\mathrm{Spec}\left(F^{0}[x_{0},\dotsc,x_{d}]/\langle x_{0}\cdots x_{a}-\lambda\rangle\right)$ as in Definition 2.11, such that $S\cap U$ is given by $\varphi^{*}(x_{0})=\cdots=\varphi^{*}(x_{a+b})=0$ for some $0\leq b\leq d-a$. Then the tropicalization
$\mathrm{trop}(\eta)=\left(\nu\left(\varphi^{*}(x_{0})(\eta)\right),\dotsc,\nu\left(\varphi^{*}(x_{a+b})(\eta)\right)\right)$
is contained in the interior of $\Delta_{S}=\Delta(a,\lambda)\times\mathbb{R}^{b}_{\geq 0}$. Without loss of generality we may assume that $\lambda=\pi^{l}$, where $l=\nu(\lambda)$ is the length of $S$. We denote by $\widetilde{\mathcal{C}}_{S}$ and $\widetilde{\mathcal{C}}_{\overline{S}}$ the restriction of $\widetilde{\mathcal{C}}$ to $S$ and $\overline{S}$, respectively. Similarly, for an irreducible component $D_{i}\subseteq D_{B^{0}}$, given locally by $\varphi^{*}(x_{i})=0$ for some $0\leq i\leq a+b$, we denote by ${\mathcal{C}}_{D_{i}}$ the restriction of ${\mathcal{C}}^{0}$ to $D_{i}$. Finally, we denote by $\psi$ the map from the preimage of $U$ in ${\mathcal{C}}^{0}$ to $\mathrm{Spec}\left(F^{0}[x_{0},\dotsc,x_{d}]/\langle x_{0}\cdots x_{a}-\pi^{l}\rangle\right)$.
3.1. Proof of part (1) Theorem 3.1
We need to specify a datum (†) as in Section 2.2.3 that satisfies the compatibility conditions of Definition 2.7. Since the tropicalizations of $K$-points give rise only to rational points of $\Lambda$, we will work with those and extend the family by linearity to the non-rational points in the very end. To ease the notation, for a face $W=\Delta_{T}\in{\mathcal{F}}(\Lambda)$ corresponding to a stratum $T\in\mathrm{Str}(B^{0},H^{0})$, set $\Theta_{T}:=\Theta_{W}$, ${\mathbb{G}}_{T}:={\mathbb{G}}_{W}$, $\ell_{T}:=\ell_{W}$, and $h_{T}:=h_{W}$. Similarly, if $W^{\prime}=\Delta_{T^{\prime}}\in{\mathcal{F}}(\Lambda)$ is such that $W$ is a face of $W^{\prime}$, i.e., $T^{\prime}\leq T$, then set $\phi_{T^{\prime},T}:=\phi_{W^{\prime},W}$.
Step 1: The tropicalization $\mathrm{trop}({\mathcal{C}}_{\eta})$ depends only on $\mathrm{trop}(\eta)\in\Lambda$, the underlying graph of $\mathrm{trop}({\mathcal{C}}_{\eta})$ depends only on $S$, and the length of each edge of $\mathrm{trop}({\mathcal{C}}_{\eta})$ is an integral affine function over $\Delta_{S}^{\circ}$ which extends to $\Delta_{S}$ in a compatible way.
By definition, the underlying graph of $\mathrm{trop}({\mathcal{C}}_{\eta})$ is the dual graph $\mathbb{G}_{s}$ of $\widetilde{\mathcal{C}}_{s}$. We first show that $\mathbb{G}_{s}$ only depends on the stratum $S$. Let $z$ be a node of $\widetilde{\mathcal{C}}_{s}$. Since ${\mathcal{C}}^{0}$ is pulled back from $\overline{M}$, there exist an étale neighborhood $V$ of $s$ in $B^{0}$ and a function $g_{z}\in\mathcal{O}_{V}(V)$ vanishing at $s$ such that the family ${\mathcal{C}}^{0}\times_{B^{0}}V$ is given by $xy=g_{z}$tale locally near $z$. After shrinking $V$, we may assume that the latter is true in a neighborhood of any node of $\widetilde{\mathcal{C}}_{s}$, and that the morphism $V\to B^{0}$ factors through $U\subseteq B^{0}$. By assumption, ${\mathcal{C}}$ is smooth over $B\backslash H$, hence each $g_{z}$ is invertible on the complement of $D_{B^{0}}$. However, $V$ is normal, and $g_{z}(s)=0$ for all nodes $z$, hence each $g_{z}$ must vanish along some component of $D_{B^{0}}\cap V$ that intersects $S$, and therefore contains $V\cap S$. Thus, all the $g_{z}$’s vanish identically along $S$. This implies that the dual graph $\mathbb{G}_{s}$ is étale locally constant along $S$. Moreover, since $\widetilde{\mathcal{C}}_{S}\rightarrow S$ is split, the identification of $\mathbb{G}_{s}$ with the dual graph of the fiber over the generic point of $S$ is canonical. Therefore, $\mathbb{G}_{s}$ only depends on $S$. Set ${\mathbb{G}}_{S}:={\mathbb{G}}_{s}$. Then for any $S\leq T\in\mathrm{Str}(B^{0},H^{0})$ there is a natural contraction $\phi_{S,T}\colon{\mathbb{G}}_{S}\rightarrow{\mathbb{G}}_{T}$ obtained by the specialization of the curve over the generic point of $T$ to the curve over the generic point of $S$.
Pick a node $z\in\widetilde{{\mathcal{C}}}_{s}$. We claim that there exist integers $n_{i}$ and $n_{\pi}$ such that
$g_{z}\cdot\pi^{-n_{\pi}}\cdot\prod_{i=0}^{a+b}\varphi^{*}(x_{i})^{-n_{i}}$ (3.1)
is regular and invertible along $S$. Here, by abuse of notation, we denote by $\varphi^{*}(x_{i})$ the pullback of the monomial $x_{i}$ to $V$ rather than $U$. Consider the log-structure on ${\rm Spec}(F^{0})$ given by the uniformizer $\pi$, the monomial log-structure on $\mathrm{Spec}(F^{0}[x_{0},\dotsc,x_{d}]/\langle x_{0}\cdot\ldots\cdot x_{a}-\pi^{l}\rangle)$ given by $\pi$ and the monomials $x_{i}$’s, and the monomial log-structure on $V$ given by the pullback of the monomial log-structure on $\mathrm{Spec}(F^{0}[x_{0},\dotsc,x_{d}]/\langle x_{0}\cdot\ldots\cdot x_{a}-\pi^{l}\rangle)$. Then
$\mathrm{Spec}(F^{0}[x_{0},\dotsc,x_{d}]/\langle x_{0}\cdot\ldots\cdot x_{a}-\pi^{l}\rangle)\to{\rm Spec}(F^{0})$
is log-smooth by [Kat94, §8.1] and $V\to\mathrm{Spec}(F^{0}[x_{0},\dotsc,x_{d}]/\langle x_{0}\cdot\ldots\cdot x_{a}-\pi^{l}\rangle)$ log-étale by [Kat89, Proposition 3.8]. Since ${\rm Spec}(F^{0})$ is log-regular, so is $V$ by [Kat94, §8.2]. Now, since $V$ is log-regular, the monomial log-structure is given by the pushforward of the sheaf of invertible functions on the complement of $D_{B^{0}}$ by [Kat94, Theorem 11.6]. Finally, since $g_{z}$ is invertible away from $D_{B^{0}}$, it follows that the function $g_{z}$ is a monomial in $\pi$ and $\varphi^{*}(x_{i})$’s up-to an invertible function as asserted. Proofs of similar claims in slightly different settings can be found in [Wło22, Lemma 4.8.2] and [Gub07, Proposition 2.11].
Next, we compute the lengths of the edges of ${\mathbb{G}}_{s}$ in $\mathrm{trop}({\mathcal{C}}_{\eta})$. By (3.1), the length of $\gamma\in E(\mathbb{G}_{s})$ corresponding to $z$ is given by
$\ell(\gamma)=\nu(g_{z}(\eta))=n_{\pi}+\sum_{i=0}^{a+b}n_{i}\cdot\nu\left(\varphi^{*}(x_{i})(\eta)\right)=\sum_{i=0}^{a+b}k_{i}\cdot\nu\left(\varphi^{*}(x_{i})(\eta)\right),$
where $k_{i}:=n_{i}+\frac{1}{l}n_{\pi}=\frac{1}{l}\mathrm{ord}_{D_{i}}(g_{z})$ if $i\leq a$ and $k_{i}:=n_{i}=\mathrm{ord}_{D_{i}}(g_{z})$ otherwise. We claim that the $k_{i}$’s are independent of the choice of $s\in S$. Indeed, for another $s^{\prime}\in S$, the corresponding $g_{z}$ and $g^{\prime}_{z}$ differ by an invertible function on the intersection of the corresponding étale neighborhoods of $s$ and $s^{\prime}$, and hence also at the generic points of the $D_{i}$’s for $0\leq i\leq a+b$. Thus, $\mathrm{ord}_{D_{i}}(g_{z})=\mathrm{ord}_{D_{i}}(g^{\prime}_{z})$ and therefore $k_{i}=k^{\prime}_{i}$ as claimed. Furthermore, we see that the length function $\ell(\gamma)$ depends only on the tropicalization $q:=\mathrm{trop}(\eta)=\left(\nu\left(\varphi^{*}(x_{0})(\eta)\right),\dotsc,\nu\left(\varphi^{*}(x_{a+b})(\eta)\right)\right)$ rather than the point $\eta$ itself.
Set $y_{i}:=\nu\left(\varphi^{*}(x_{i})(\eta)\right)$. Then, $q=(y_{0},\dotsc,y_{a+b})$ and the length function
$\ell_{S}(\gamma,q)=n_{\pi}+\sum_{i=0}^{a+b}n_{i}y_{i}=\sum_{i=0}^{a+b}k_{i}y_{i}$
is an integral affine function on $\Delta_{S}^{\circ}$ that extends naturally to $\Delta_{S}$. It remains to show that $\ell_{S}(\gamma,q)$ satisfies the compatibility condition of Definition 2.7 (2). Note that $g_{z}$ is defined over an étale open subset of $B^{0}$ which intersects all strata $T\in\mathrm{Str}(B^{0},H^{0})$ satisfying $S\leq T$. Let $T$ be the stratum given locally by
$\varphi^{*}(x_{0})=\cdots=\varphi^{*}(x_{a^{\prime}})=0=\varphi^{*}(x_{a+1})=\cdots=\varphi^{*}(x_{a+b^{\prime}}),$
for some $a^{\prime}\leq a$ and $b^{\prime}\leq b$. Then $\Delta_{T}$ is the face of $\Delta_{S}$ given by
$y_{a^{\prime}+1}=\cdots=y_{a}=y_{a+b^{\prime}+1}=\cdots=y_{a+b}=0.$
If the node $z$ is smoothed out in the family of curves over $T$, then $g_{z}$ is invertible at the generic point of $D_{i}$ with $0\leq i\leq a^{\prime}$ or $a+1\leq i\leq a+b^{\prime}$, and the edge $\gamma$ gets contracted to a vertex in ${\mathbb{G}}_{T}$. Hence $k_{i}=0$ for $0\leq i\leq a^{\prime}$ or $a+1\leq i\leq a+b^{\prime}$ and the compatibility holds:
$\ell_{T}(\phi_{S,T}(\gamma),q)=0=\ell_{S}(\gamma,q),\ \mathrm{for\ all}\ q\in\Delta_{T}.$
Assume now that $z$ is the specialization of a node of the curve over the generic point of $T$. Then $\gamma$ is not contracted by $\phi_{S,T}$. In this case, again for any $s^{\prime}\in T$, the corresponding $g^{\prime}_{z}$ and $g_{z}$ differ by a function invertible at the generic points of all $D_{i}$ that contain $T$, namely, with $0\leq i\leq a^{\prime}$ or $a+1\leq i\leq a+b^{\prime}$. Therefore, $\mathrm{ord}_{D_{i}}(g_{z})=\mathrm{ord}_{D_{i}}(g^{\prime}_{z})$ for those $D_{i}$’s, and hence the restriction of $\ell_{S}(\gamma,\cdot)$ to $\Delta_{T}$ coincides with $\ell_{T}(\phi_{S,T}(\gamma),\cdot)$ as needed.
Step 2: The parameterization $h_{\eta}\colon\mathrm{trop}({\mathcal{C}}_{\eta})\rightarrow N_{\mathbb{R}}$ depends only on $\mathrm{trop}(\eta)$, and varies integral affine linearly on $\Delta_{S}^{\circ}$ with constant combinatorial type. It also extends to $\Delta_{S}$ in a compatible way.
Since each section $\sigma^{0}_{\bullet}$ of the family ${\mathcal{C}}\rightarrow B^{0}$ is mapped to a unique toric divisor or the dense orbit in $X$, the slopes of the legs of $\mathrm{trop}({\mathcal{C}}_{\eta})$ are independent of $\eta$ by Remark 2.10. Therefore, the extended degree of $h_{\eta}\colon\mathrm{trop}({\mathcal{C}}_{\eta})\rightarrow N_{\mathbb{R}}$ is independent of $\eta$ and will be denoted by $\overline{\nabla}$.
Recall that the closure $\overline{S}$ of $S$ is a component of the intersection of $D_{0},\dotsc,D_{a+b}$. Let $u$ be a vertex of ${\mathbb{G}}_{S}$. Denote by $\widetilde{\mathcal{C}}_{u}$ the component of $\widetilde{\mathcal{C}}_{S}$ corresponding to $u$ and by $\widetilde{\mathcal{C}}_{s,u}$ the component of $\widetilde{\mathcal{C}}_{s}$ corresponding to $u$. As in Step 1, let $q=(y_{0},\dotsc,y_{a+b}):=\mathrm{trop}(\eta)$. Let $\widetilde{\mathcal{C}}_{u}^{\prime}\subset\widetilde{\mathcal{C}}_{u}$ be the open subset of non-special points. Then in a neighborhood of $\widetilde{\mathcal{C}}_{u}^{\prime}$, the toroidal structure on $\mathscr{C}^{0}$ is pulled back from the one on $B^{0}$. Therefore, as in Step 1, for any $m\in M$, there exist integers $n_{i}$ for $0\leq i\leq a+b$ and $n_{\pi}$ such that
$f^{*}(x^{m})\cdot\pi^{-n_{\pi}}\cdot\prod_{0\leq i\leq a+b}\psi^{*}(x_{i})^{-n_{i}}$
is regular and invertible along $\widetilde{\mathcal{C}}_{u}^{\prime}$. Hence, by definition, the inner product of $h_{\eta}(u)\in N_{\mathbb{R}}$ with $m$ is given by
$-\mathrm{ord}_{\widetilde{\mathcal{C}}_{s,u}}\left(f^{*}(x^{m})|_{{\mathcal{C}}_{\overline{\eta}}}\right)=-n_{\pi}-\sum_{i=0}^{a+b}n_{i}\cdot\nu\left(\varphi^{*}(x_{i})(\eta)\right)=-n_{\pi}-\sum_{i=0}^{a+b}n_{i}y_{i},$
which only depends on $q=\mathrm{trop}(\eta).$ Let us write $h_{S}(u,q)=h_{\eta}(u)$, then $h_{S}(u,\cdot)$ is an integral affine function on $\Delta_{S}$. By a similar argument as in Step 1, we see that this function satisfies the compatibility of Definition 2.7 (3).
As the parameterization $h_{\eta}$ varies continuously with respect to $q=\mathrm{trop}(\eta)$, while the set of combinatorial types with underlying graph ${\mathbb{G}}_{S}$ is a discrete set, we see that the combinatorial type of $h_{\eta}$ is constant, which we denote by $\Theta_{S}$.
Step 3: The conclusion of the proof of (1).
In the first two steps, we have constructed the data of a family of parameterized topical curve over $\Lambda$ the extended degree, the combinatorial type for each face of $\Lambda$, the weighted contractions, and the integral affine functions $\ell_{S}$ and $h_{S}$. Furthermore, we have verified the compatibility conditions of Definition 2.7, and have seen that the resulting family of parameterized topical curves over $\Lambda$ satisfies assertion (1).
Proof of Part (2) of Theorem 3.1
Denote by $\chi\colon B^{0}\rightarrow\overline{M}$ the map inducing the family ${\mathcal{C}}^{0}$. Let $S\in\mathrm{Str}(B^{0},H^{0})$ be the stratum corresponding to $W$. Then the closure $\overline{S}$ of $S$ is a proper and smooth variety. Denote by $\Theta$ the combinatorial type of $\alpha(W^{\circ})$. There are two cases to consider:
Case 1: $\chi$ contracts $\overline{S}$. In this case we will prove that $\alpha$ is harmonic along $W$ if $W$ has codimension one and quasi-harmonic along $W$ otherwise. Set $p:=\chi(\overline{S})$ and let $C_{p}$ be the fiber of the universal family over $p$. Since ${\mathcal{C}}^{0}$ is a pullback of the universal family along $\chi$, it follows that $\widetilde{\mathcal{C}}_{\overline{S}}=\overline{S}\times C_{p}$. Furthermore, ${\mathbb{G}}_{S}$ is the dual graph of $C_{p}$, and $\alpha$ maps ${\rm Star}(W^{\circ})$ to $M_{[\Theta]}$, which lifts to a map to $M_{\Theta}\subset{\mathbb{R}}^{|E({\mathbb{G}}_{S})|}\times N_{\mathbb{R}}^{|V({\mathbb{G}}_{S})|}$. By abuse of notation, the lifting is also denoted by $\alpha$.
Recall that each point $s\in S$ admits an open neighborhood $s\in U\subset B^{0}$ and an étale morphism $\varphi\colon U\rightarrow\mathrm{Spec}\left(F^{0}[x_{0},\dotsc,x_{d}]/\langle x_{0}\cdots x_{a}-\pi^{l}\rangle\right)$ such that $S\cap U$ is given by $\varphi^{*}(x_{0})=\cdots=\varphi^{*}(x_{a+b})=0$ for some $0\leq b\leq d-a$, and hence $W=\Delta(a,\pi^{l})\times\mathbb{R}^{b}_{\geq 0}$. Let $v_{0},\dotsc,v_{a}$ be the vertices of $W$. For $a+1\leq i\leq a+b$, let $\vec{e}_{i}$ be the standard basis of $\mathbb{R}^{b}$ and set $v_{i}=v_{0}+\vec{e}_{i}$. Then ${\rm Lin}(\alpha(W))$ is generated by $\alpha(v_{i})-\alpha(v_{0})$ for $1\leq i\leq a+b$.
Let $\{S_{i}\}$, where $a+b+1\leq i\leq a+b+r+t$, be the set of strata of codimension one in $\overline{S}$. Each $S_{i}$ corresponds to a face $W_{i}\in\mathcal{F}(\Lambda)$ containing $W$ as a codimension-one face. Suppose $\overline{S}_{i}=\overline{S}\cap D_{i}$ for some component $D_{i}$ of $D_{B^{0}}$. We may further assume that $D_{i}$ is vertical for $a+b+1\leq i\leq a+b+r$, and horizontal for $a+b+r+1\leq i\leq a+b+r+t$. Note that the $D_{i}$’s are uniquely determined by the $S_{i}$’s but may not necessarily be distinct. In the sequel, we will mainly consider the following components of $D_{B^{0}}$, which have non-empty intersection with $\overline{S}$: the components $D_{0},\dotsc,D_{a+b}$, which are distinct and contain $\overline{S}$; and the components $D_{a+b+1},\dotsc,D_{a+b+r+t}$ constructed above.
For $a+b+1\leq i\leq a+b+r$, let $l_{i}$ be the length of $S_{i}$ and $v^{\prime}_{i}$ be the vertex of $W_{i}$ not contained in $W$. In fact, $l_{i}=l$ unless $a=0$. Let $\vec{e}_{i}=\frac{1}{l_{i}}(v^{\prime}_{i}-v_{0})$ be the primitive integral vector parallel to $v^{\prime}_{i}-v_{0}$, and set $v_{i}:=v_{0}+\vec{e}_{i}$. For $a+b+r+1\leq i\leq a+b+r+t$ we have $W_{i}=W\times{\mathbb{R}}_{\geq 0}$. In this case, set $v_{i}:=v_{0}+\vec{e}_{i}$, where $\vec{e}_{i}$ is the unit normal vector to $W$ in $W_{i}$. Then $W_{i}$ is contained in $W+{\mathbb{R}}_{\geq 0}\vec{e}_{i}$ for $a+b+1\leq i\leq a+b+r+t$. Consequently, we have ${\rm Star}(W)=\{\vec{e}_{i}+N_{W}\}_{a+b+1\leq i\leq a+b+r+t}$ and $\frac{\partial\alpha}{\partial\vec{e}_{i}}=\alpha(v_{i})-\alpha(v_{0})$. Thus, to prove quasi-harmonicity (resp. harmonicity), it remains to show that
$\sum_{i=a+b+1}^{a+b+r+t}a_{i}\left(\alpha(v_{i})-\alpha(v_{0})\right)\in{\rm Lin}(\alpha(W))$ (3.2)
for some positive integers $a_{i}$ (resp. $a_{i}=1$).
Since $\overline{S}$ is étale locally isomorphic to an affine space and the $\overline{S}_{i}$’s correspond to the coordinate hyperplanes, we can pick a (possibly reducible) curve $C$ in $\overline{S}$ that has positive intersection numbers with each $\overline{S}_{i}$ and none of whose components is contained in $\cup\overline{S}_{i}$. In particular, when $\dim S=1$ (equivalently, $W$ has codimension one in $\Lambda$), we set $C:=\overline{S}$. Set $l_{i}:=l$ for $i\leq a$, and $l_{i}:=1$ for $a+1\leq i\leq a+b$ and for $i>a+b+r$. Then (3.2) would follow from the following equality
$\sum_{i=a+b+1}^{a+b+r+t}\overline{S}_{i}\cdot C\ \big{(}\alpha(v_{i})-\alpha(v_{0})\big{)}=-\sum_{i=0}^{a+b}l_{i}D_{i}\cdot C\ \big{(}\alpha(v_{i})-\alpha(v_{0})\big{)},$ (3.3)
where the intersection product on the left is in $\overline{S}$, while the one on the right is in $B^{0}$. Note that when $\dim S=1$ we have $\overline{S}_{i}\cdot C=1$ by construction.
We now prove (3.3) coordinatewise. Let $\gamma\in E({\mathbb{G}}_{S})$ be an edge corresponding to a node $z\in C_{p}$. Let $u,u^{\prime}\in V({\mathbb{G}}_{S})$ be the vertices adjacent to $\gamma$, let $C_{u}$ and $C_{u^{\prime}}$ be the corresponding components of $C_{p}$, and $\widetilde{{\mathcal{C}}}_{u},\widetilde{{\mathcal{C}}}_{u^{\prime}},Z$ the pullbacks of $C_{u},C_{u^{\prime}},z$ to $\widetilde{{\mathcal{C}}}_{\overline{S}}$, respectively. The universal curve over $\overline{M}$ is given étale locally at $z$ by $xy=m_{z}$, where $m_{z}$ is defined on an étale neighborhood of $p$, and vanishes at $p$. Thus, ${\mathcal{C}}^{0}$ is given by $xy=g_{Z}:=\chi^{*}(m_{z})$ in an étale neighborhood of $Z$. Notice that we constructed a function defined in a neighborhood of the whole family of nodes $Z$, which globalizes the local construction of Step 1 in the particular case we consider here.
According to Step 1, the $\gamma$-coordinate of $\alpha(v_{i})-\alpha(v_{0})$ is equal to $\frac{1}{l_{i}}\left({\rm ord}_{D_{i}}(g_{Z})-{\rm ord}_{D_{0}}(g_{Z})\right)$ when $D_{i}$ is vertical, namely $0\leq i\leq a$ or $a+b+1\leq i\leq a+b+r$; and equal to ${\rm ord}_{D_{i}}(g_{Z})$ otherwise. Set $k_{i}:=\frac{1}{l_{i}}{\rm ord}_{D_{i}}(g_{Z})$.
Then the $\gamma$-coordinate of the left-hand side of (3.3) is given by
$\sum_{i=a+b+1}^{a+b+r+t}k_{i}\overline{S}_{i}\cdot C-\sum_{i=a+b+1}^{a+b+r}k_{0}\frac{l}{l_{i}}\overline{S}_{i}\cdot C;$
while on the right-hand side it is given by
$-\sum_{i=0}^{a+b}l_{i}k_{i}D_{i}\cdot C+\sum_{i=0}^{a}l_{0}k_{0}D_{i}\cdot C=-\sum_{i=0}^{a+b}l_{i}k_{i}D_{i}\cdot C+\sum_{i=0}^{a}lk_{0}D_{i}\cdot C.$
Therefore, (3.3) is equivalent to
$\sum_{i=0}^{a+b}l_{i}k_{i}D_{i}\cdot C+\sum_{i=a+b+1}^{a+b+r+t}k_{i}\overline{S}_{i}\cdot C=lk_{0}\Big{(}\sum_{i=0}^{a}D_{i}\cdot C+\sum_{i=a+b+1}^{a+b+r}\frac{1}{l_{i}}\overline{S}_{i}\cdot C\Big{)},$ (3.4)
which holds true since both sides vanish. Indeed, since $\overline{S}_{i}\cdot C=l_{i}D_{i}\cdot C$ for all $i>a+b$, the left-hand side of (3.4) is $\mathrm{div}(g_{Z})\cdot C$, which vanishes since $C$ is complete, while the right-hand side is $k_{0}l\widetilde{B}\cdot C=0$ since $\widetilde{B}$ is a principal divisor of $B^{0}$ defined by $\pi$.
Next we verify (3.3) for the $u$-coordinates. Pick $u\in V({\mathbb{G}}_{S})$. For $1\leq i\leq a+b+r+t$, let ${\mathcal{C}}_{i}$ be the irreducible component of ${\mathcal{C}}_{D_{i}}$ that contains some fiber of $\widetilde{\mathcal{C}}_{u}$. Denote $k^{\prime}_{i}=\frac{1}{l_{i}}{\rm ord}_{{\mathcal{C}}_{i}}f^{*}(x^{m})$. In this case, by taking inner product of both sides of (3.3) with an arbitrary $m\in M$, a similar computation as above shows that (3.3) is equivalent to
$\sum_{i=0}^{a+b}l_{i}k^{\prime}_{i}D_{i}\cdot C+\sum_{i=a+b+1}^{a+b+r+t}k^{\prime}_{i}\overline{S}_{i}\cdot C=lk^{\prime}_{0}\Big{(}\sum_{i=0}^{a}D_{i}\cdot C+\sum_{i=a+b+1}^{a+b+r}\frac{1}{l_{i}}\overline{S}_{i}\cdot C\Big{)}.$
Take a section $\widetilde{\sigma}_{S}$ of $\widetilde{\mathcal{C}}_{u}\rightarrow\overline{S}$ that is disjoint with any section $\sigma^{0}_{\bullet}$ of ${\mathcal{C}}^{0}$ and the singular locus of ${\mathcal{C}}^{0}$, and let $\widetilde{\sigma}_{C}$ (resp. $\widetilde{\sigma}_{i}$) be the restriction of $\widetilde{\sigma}_{S}$ on $C$ (resp. $\overline{S}_{i}$). By projection formula, we have $D_{i}\cdot C={\mathcal{C}}_{i}\cdot\widetilde{\sigma}_{C}$. Then (3.3) is equivalent to
$\sum_{i=0}^{a+b}l_{i}k^{\prime}_{i}\mathscr{C}_{i}\cdot\widetilde{\sigma}_{C}+\sum_{i=a+b+1}^{a+b+r+t}k^{\prime}_{i}\widetilde{\sigma}_{i}\cdot\widetilde{\sigma}_{C}=lk^{\prime}_{0}\Big{(}\sum_{i=0}^{a}D_{i}\cdot C+\sum_{i=a+b+1}^{a+b+r}\frac{1}{l_{i}}\overline{S}_{i}\cdot C\Big{)},$
and again, the latter equality holds true since both sides vanish. Indeed, the left-hand side is now the intersection $\widetilde{\sigma}_{C}\cdot\mathrm{div}f^{*}(x^{m})$ and the right-hand side is $lk^{\prime}_{0}\widetilde{B}\cdot C$. This concludes the proof of (3.3) and of Case 1.
Case 2: $\chi$ does not contract $\overline{S}$. By the assumptions of the theorem, the graph ${\mathbb{G}}_{S}$ is weightless and $3$-valent except for at most one $4$-valent vertex. Since rational curves with three marked points have no moduli, it follows that ${\mathbb{G}}_{S}$ has a $4$-valent vertex, which we denote by $u\in{\mathbb{G}}_{S}$. In this case we will show that the map $\alpha$ is locally combinatorially surjective at $W$. Consider $\widetilde{\mathcal{C}}_{u}$ as above. By construction, $\widetilde{\mathcal{C}}_{u}\to\overline{S}$ is a family of rational curves and since $\widetilde{\mathcal{C}}\to\widetilde{B}$ is split, the four marked points in each fiber define sections of the family. Let $\xi\colon\overline{S}\to\overline{M}_{0,4}\simeq{\mathbb{P}}^{1}$ be the induced map. Since $\chi$ does not contract $\overline{S}$, it follows that $\xi$ is not constant, and hence surjective. Thus, the preimage of each boundary point of $\overline{M}_{0,4}$ in $\overline{S}$ is the union of closures of some codimension-one strata $S^{\prime}$. Such strata correspond to faces $W^{\prime}\in\mathcal{F}(\Lambda)$ containing $W$ as a codimension-one face. Since $\xi$ is surjective, it follows that for each weightless and $3$-valent types $\Theta^{\prime}$ corresponding to one of the three possible resolutions of the $4$-valent vertex $u$, there exists a face $W^{\prime}$ as above for which $\alpha(W^{\prime})\cap M_{[\Theta^{\prime}]}\neq\emptyset$. Therefore $\alpha$ is locally combinatorially surjective along $W$ as claimed. ∎